Draw the graph of these eqns.  And explain also how?


X2+2x+1=0


-X2-2x-1=0


 



Grade:12th Pass

1 Answers

Vinayak Gupta
12 years ago

Both the mentioned equations are same and the solution is x = -1 .

multiplying equation 2 with -1 we will get the first equation.

the equation 1 is in the form of  

(a+b)2 = a2 + b2 + 2ab

we will get ,

(x+1) = 0.

hence x = -1.

the graph of the equation will be a straight line passing through the x axis at -1.

if you want i will paste the graph also.

please approve the answer if helpful .